# Conveyor: Efficient Tool-aware LLM Serving with Tool Partial Execution

### v1

@@ 1 Introduction
The rapid evolution of large language models (LLMs) has significantly accelerated in recent years, and LLMs have quickly become the state-of-the-art approach in many AI tasks, such as content generation, question answering, and text classification. Consequently, LLM serving systems have emerged as a crucial component in deploying these models in various applications to achieve high performance and resource efficiency. Many LLM serving techniques have been proposed to reduce response latency (18; 4; 9; 3) and improve system throughput (22; 16; 28; 7; 6; 5).
Recently, a new use case, tool-assisted LLM serving, has emerged to enhance the reasoning capabilities of LLMs and enable them to interact with the external world. In a typical tool-assisted LLM serving workflow, a user first sends a request (i.e., the original prompt) to the system. An LLM will then process this request and generate a set of tool calling commands, or plans. Tool executors will invoke various tools to execute these commands and collect outputs, or observations. The original prompt, plans, and observations will be concatenated following a template, and then sent back to the LLM. The LLM generates either new plans, indicating a new round of tool execution, or the ultimate response to be sent back to the user. Example tools include but are not limited to calculators, databases, code interpreters, search engines, and ticket-booking travel agency websites. The most notable example is ChatGPT plugins, which have allowed users to book air tickets (13) and to reason about mathematical expressions (26).
@@
We evaluate Conveyor with six LLM serving workloads that contain external tool invocations. We use Mistral-7B-Instruct-v0.2 (20) and Functionary-Small-v2.2 (10) to invoke tools. Our results show that Conveyor’s performance improvement varies significantly across different tools. Specifically, Conveyor achieves a latency improvement of up to 38.8% across workloads including code generation, search, and planning. However, Conveyor shows negligible improvement in database execution and calculator tools. We also provide mathematical analysis of when tool partial execution can bring performance improvement and show that our empirical evaluation matches the analysis. We would like to highlight that the effectiveness of Conveyor is not affected by the choice of LLM or prompts because the execution flow of LLM decoding instructions and invoking tools remains the same that tool partial execution can happen alongside LLM decoding. Our source code will be publicly available at https://github.com/conveyor-sys/conveyor.
@@ 2.1 LLM Serving Systems
Modern LLMs predominantly employ the Transformer architecture (25), at the core of which lies the self-attention module. The self-attention module computes three vectors for each token in the sequence, including query (Q), key (K), and value (V) vectors. It then calculates the attention score for each token by multiplying its Q vector with the K vectors of all preceding tokens, followed by a softmax and weighted average computation. Since the key and value vectors for processed tokens will be reused when generating new tokens, previous keys and values are usually cached in the GPU memory, known as the KV cache.
To process an LLM serving request, a Transformer-based LLM operates through two phases: Prefilling and Decoding. During the prefilling phase, the entire user prompt is processed and the LLM generates the first output token. The user prompt can be processed in parallel in a single iteration. Therefore, GPU utilization is typically high during this prefilling phase thanks to the intra-request parallelism. During the decoding phase, the model generates output tokens sequentially, relying on all previously generated tokens, including both user prompts and all tokens produced thus far. This sequential generation inherently results in lower GPU utilization and throughput as only one token can be produced per iteration for a single serving request. This sequential generation process is called autoregressive decoding. Therefore, modern LLM serving systems batch multiple serving requests together to improve system throughput and resource utilization. For example, continuous batching (28; 16) has become the most widely deployed batching technique in existing LLM serving systems.
@@ 2.2 LLM Serving with External Tool Invocations
Recently, there has been a rising trend of integrating LLM serving with external tool execution (e.g., ChatGPT plugins (13; 26) and Toolformer (24)). External tools extend an LLM’s capabilities to perform various complex tasks and interact with external environments. For example, many AI agents enable LLMs to access search engines to acquire up-to-date information or access databases that contain private datasets. LLMs can also invoke calculators to reason about complex math equations and execute code interpreters to generate customized data visualization. Moreover, they can trigger other ML models (e.g., computer vision models) to understand image contents.
@@
Integrating external tools into an LLM serving system has inspired a new line in machine learning system research: KV cache management. Since multiple rounds of LLM serving are typically needed for a single user request, the resources (e.g., KV cache) for a finished LLM serving are likely to be reused by future LLM inferences (1). Treating these multiple rounds of LLM serving as independent requests results in redundant prefilling of the same token sequences. For example, in Figure 2, to process the serving request (7), the LLM needs to compute the KV vectors for the original prompt and generated plans, which have already been computed previously as (2) and (3). AttentionStore (11) evicts the KV cache to CPU memory as long as the PCIe bandwidth permits the transfer. When the next round initiates, the KV cache is then loaded back into the GPU memory to eliminate KV cache recomputation. APIServe (1) predicts the execution time of external API access and estimates the corresponding GPU resource waste. APIServe then uses the estimation results to make GPU management decisions, such as discarding the key-value states, swapping the states to CPU memory, or retaining the states inside the GPU. Managing the KV cache is only one of the opportunities to accelerate LLM serving with external tools shown in this of works. Next, we will show that there is one additional opportunity to improve LLM serving with external tool invocations.
@@ 3.1 Efficiency Opportunities in Modern Tool-assisted LLM Serving Systems
Today’s tool-assisted LLM serving systems start the plan execution only after the LLM completes the entire decoding procedure. This misses the opportunities for pipelining LLM’s decoding and plan execution to achieve reduced serving latency and improved system efficiency. For example, it would be ideal if the Python interpreter could start partially executing the script as soon as the LLM generates the first line of code (e.g., import torch), without waiting for the generation of the entire script. However, in existing designs (17; 24; 1), since the LLM serving system and the tool execution are not co-optimized, the Python interpreter will only start after the entire script has been generated. This leads to extended serving delay and inefficient resource utilization (e.g., the LLM is idle and GPU cycles are wasted). We name this desired capability of initiating tool execution before complete LLM decoding as tool partial execution, which has the potential to significantly reduce the serving latency and improve the overall efficiency and responsiveness of the system.
{minted}
[xleftmargin=20pt,linenos,fontsize=]python “‘python import torch import torchvision.utils as vutils import matplotlib.pyplot as plt import numpy as np
x = torch.linspace(0, 2 * torch.tensor(np.pi), 1000) y = torch.sin(x)
vutils.save_image(y.unsqueeze(0), ’sin_wave.png’) plt.plot(x.numpy(), y.numpy()) plt.show() “‘
@@
Let’s consider a concrete example of executing Python code (in Figure 3). This code is generated by Mistral-7B-Instruct-v0.2 (20) with the prompt “Plotting a sine wave in python with torch and matplotlib. ONLY output code without trailing explanation.”. We use the markdown code block syntax ```python and ``` as the indicators for the start and end of the tool. The execution without tool partial execution is that the LLM first generates the entire Python script, executes the script, and returns the image to the user. To understand why Conveyor provides performance improvement in this case, we plot the execution timeline with and without tool partial execution in Figure 4. The green blocks represent the LLM decoding of a line of Python. The numbers represent line numbers. The grey boxes represent the execution of the Python code in a Python interpreter. With partial execution, the execution of lines 1–12 can be completely pipelined with the decoding procedure, and only line 13 needs to be executed after the decoding is finished.
@@ 3.3 Conveyor Execution Workflow
It is worthwhile to note that we choose to spawn tool execution on a separate dedicated process to avoid unnecessary contention or interference with the LLM decoding procedure. Machine learning serving software usually uses Python as the programming language platform (e.g., PyTorch (21)), and running a Python-based tool in the same process may cause either the tool execution or LLM decoding to be blocked by the Python Global Interpreter Lock (GIL) mechanism. Such contention can cause extended latency and reduce system performance. Another feature of Conveyor is that the parser implementation only depends on the syntax of the tool message generated by the model. Therefore, the number of parsers typically only depends on the number of models supported by the system, which is advantageous for tool developers.
@@ 4.1 Setup
We evaluate Conveyor on our testbed of servers with two Intel 10-core Xeon Gold 5215 CPUs (running at 2.5 GHz base frequency) and two NVIDIA GeForce RTX 3090 GPUs. We implement our system on top of PyTorch (21) with FlashInfer CUDA kernels (8). Our Conveyor system is implemented in about 1.9K lines of Python. Our baseline is the same code but with tool partial execution disabled, where tool invocation always happens after decoding to the end-of-sequence (EOS) token.
@@
To the best of our knowledge, there are unfortunately no publicly available realistic datasets of tool-assisted LLM serving scenarios. Instead, we systematically investigate existing literature (14; 19; 12; 2; 27; 23; 15) and construct six scenarios for our evaluation.
@@
Due to the models’ capabilities limited by our testbed (the available GPU memory), we generate these synthetic workloads with our own prompts instead of using the original prompts of those papers. We use Mistral-7B-Instruct-v0.2 (20) for Python CodeGen and Planning workloads, and use Functionary-Small-v2.2 (10) for Search, Validation, Database, and Calculator workloads. We choose these two models because they can generate valid tool invocation instructions for us to run the evaluation except for the Validation workload. We pick Functionary for the Validation workload because we empirically found that it has a higher probability of generating format-correct requests. We set temperature to be 0, so every test for the same workload has the same LLM output. Performance variance across tests for the same workload is due to the performance variance in CPU/GPU processing and the Internet (for the Search and Planning workloads). Note that the effectiveness of Conveyor is not affected by the choice of LLM or prompts because the execution flow of LLM decoding instructions and invoking tools remains the same, which will be further elaborated in the following case study section.
@@ 4.3 Case Studies
For the CodeGen workload, we let Mistral-7B-Instruct-v0.2 (20) generates the Python code to plot a sine wave (Figure 3). Figure 4 shows the corresponding execution timelines with and without tool partial execution. Conveyor helps to reduce averagely 725 ms for the entire end-to-end serving latency, which is 3,918 ms on average without partial execution, leading to 26.3% improvement (see § 3.4) for this user request. It is worthwhile to note that a few import statements (e.g., line 2 and line 5) consume negligible time because Conveyor invokes the Python interpreter through fork, so the plotting script can reuse the modules already imported by Conveyor.
@@
For the Validation workload, we let Functionary-Small-v2.2 (10) generate a request for a local news service. The request is in JSON format, and the field “location” has to contain a valid city name and a state name. If the “location” only contains the city name, the request will be rejected by the service, and there is no point in sending such a request. Figure 7 shows the execution timeline. Without partial execution, such a check has to take place after the entire request is decoded. Partial execution allows the check to happen earlier and can abort immediately, saving the resources and the time for decoding subsequent tokens.

### v2

@@ 1 Introduction
The rapid evolution of large language models (LLMs) has significantly accelerated in recent years, and LLMs have quickly become the state-of-the-art approach in many AI tasks, such as content generation, question answering, and text classification. Consequently, LLM serving systems have emerged as a crucial component in deploying these models in various applications to achieve high performance and resource efficiency. Many LLM serving techniques have been proposed to reduce response latency [18, 4, 9, 3] and improve system throughput [22, 16, 28, 7, 6, 5].
Recently, a new use case, tool-assisted LLM serving, has emerged to enhance the reasoning capabilities of LLMs and enable them to interact with the external world. In a typical tool-assisted LLM serving workflow, a user first sends a request (i.e., the original prompt) to the system. An LLM will then process this request and generate a set of tool calling commands, or plans. Tool executors will invoke various tools to execute these commands and collect outputs, or observations. The original prompt, plans, and observations will be concatenated following a template, and then sent back to the LLM. The LLM generates either new plans, indicating a new round of tool execution, or the ultimate response to be sent back to the user. Example tools include but are not limited to calculators, databases, code interpreters, search engines, and ticket-booking travel agency websites. The most notable example is ChatGPT plugins, which have allowed users to book air tickets [13] and to reason about mathematical expressions [26].
@@
We evaluate Conveyor with six LLM serving workloads that contain external tool invocations. We use Mistral-7B-Instruct-v0.2 [20] and Functionary-Small-v2.2 [10] to invoke tools. Our results show that Conveyor’s performance improvement varies significantly across different tools. Specifically, Conveyor achieves a latency improvement of up to 38.8% across workloads including code generation, search, and planning. However, Conveyor shows negligible improvement in database execution and calculator tools. We also provide mathematical analysis of when tool partial execution can bring performance improvement and show that our empirical evaluation matches the analysis. We would like to highlight that the effectiveness of Conveyor is not affected by the choice of LLM or prompts because the execution flow of LLM decoding instructions and invoking tools remains the same that tool partial execution can happen alongside LLM decoding. Our source code will be publicly available at https://github.com/conveyor-sys/conveyor.
@@ 2.1 LLM Serving Systems
Modern LLMs predominantly employ the Transformer architecture [25], at the core of which lies the self-attention module. The self-attention module computes three vectors for each token in the sequence, including query (Q), key (K), and value (V) vectors. It then calculates the attention score for each token by multiplying its Q vector with the K vectors of all preceding tokens, followed by a softmax and weighted average computation. Since the key and value vectors for processed tokens will be reused when generating new tokens, previous keys and values are usually cached in the GPU memory, known as the KV cache.
To process an LLM serving request, a Transformer-based LLM operates through two phases: Prefilling and Decoding. During the prefilling phase, the entire user prompt is processed and the LLM generates the first output token. The user prompt can be processed in parallel in a single iteration. Therefore, GPU utilization is typically high during this prefilling phase thanks to the intra-request parallelism. During the decoding phase, the model generates output tokens sequentially, relying on all previously generated tokens, including both user prompts and all tokens produced thus far. This sequential generation inherently results in lower GPU utilization and throughput as only one token can be produced per iteration for a single serving request. This sequential generation process is called autoregressive decoding. Therefore, modern LLM serving systems batch multiple serving requests together to improve system throughput and resource utilization. For example, continuous batching [28, 16] has become the most widely deployed batching technique in existing LLM serving systems.
@@ 2.2 LLM Serving with External Tool Invocations
Recently, there has been a rising trend of integrating LLM serving with external tool execution (e.g., ChatGPT plugins [13, 26] and Toolformer [24]). External tools extend an LLM’s capabilities to perform various complex tasks and interact with external environments. For example, many AI agents enable LLMs to access search engines to acquire up-to-date information or access databases that contain private datasets. LLMs can also invoke calculators to reason about complex math equations and execute code interpreters to generate customized data visualization. Moreover, they can trigger other ML models (e.g., computer vision models) to understand image contents.
@@
Integrating external tools into an LLM serving system has inspired a new line in machine learning system research: KV cache management. Since multiple rounds of LLM serving are typically needed for a single user request, the resources (e.g., KV cache) for a finished LLM serving are likely to be reused by future LLM inferences [1]. Treating these multiple rounds of LLM serving as independent requests results in redundant prefilling of the same token sequences. For example, in Figure 2, to process the serving request (7), the LLM needs to compute the KV vectors for the original prompt and generated plans, which have already been computed previously as (2) and (3). AttentionStore [11] evicts the KV cache to CPU memory as long as the PCIe bandwidth permits the transfer. When the next round initiates, the KV cache is then loaded back into the GPU memory to eliminate KV cache recomputation. InferCept [1] predicts the execution time of external API access and estimates the corresponding GPU resource waste. InferCept then uses the estimation results to make GPU management decisions, such as discarding the key-value states, swapping the states to CPU memory, or retaining the states inside the GPU. Managing the KV cache is only one of the opportunities to accelerate LLM serving with external tools shown in this of works. Next, we will show that there is one additional opportunity to improve LLM serving with external tool invocations.
@@ 3.1 Efficiency Opportunities in Modern Tool-assisted LLM Serving Systems
Today’s tool-assisted LLM serving systems start the plan execution only after the LLM completes the entire decoding procedure. This misses the opportunities for pipelining LLM’s decoding and plan execution to achieve reduced serving latency and improved system efficiency. For example, it would be ideal if the Python interpreter could start partially executing the script as soon as the LLM generates the first line of code (e.g., import torch), without waiting for the generation of the entire script. However, in existing designs [17, 24, 1], since the LLM serving system and the tool execution are not co-optimized, the Python interpreter will only start after the entire script has been generated. This leads to extended serving delay and inefficient resource utilization (e.g., the LLM is idle and GPU cycles are wasted). We name this desired capability of initiating tool execution before complete LLM decoding as tool partial execution, which has the potential to significantly reduce the serving latency and improve the overall efficiency and responsiveness of the system.
@@
Let’s consider a concrete example of executing Python code (in Figure 3). This code is generated by Mistral-7B-Instruct-v0.2 [20] with the prompt “Plotting a sine wave in python with torch and matplotlib. ONLY output code without trailing explanation.”. We use the markdown code block syntax ```python and ``` as the indicators for the start and end of the tool. The execution without tool partial execution is that the LLM first generates the entire Python script, executes the script, and returns the image to the user. To understand why Conveyor provides performance improvement in this case, we plot the execution timeline with and without tool partial execution in Figure 4. The green blocks represent the LLM decoding of a line of Python. The numbers represent line numbers. The grey boxes represent the execution of the Python code in a Python interpreter. With partial execution, the execution of lines 1–12 can be completely pipelined with the decoding procedure, and only line 13 needs to be executed after the decoding is finished.
@@ 3.3 Conveyor Execution Workflow
It is worthwhile to note that we choose to spawn tool execution on a separate dedicated process to avoid unnecessary contention or interference with the LLM decoding procedure. Machine learning serving software usually uses Python as the programming language platform (e.g., PyTorch [21]), and running a Python-based tool in the same process may cause either the tool execution or LLM decoding to be blocked by the Python Global Interpreter Lock (GIL) mechanism. Such contention can cause extended latency and reduce system performance. Another feature of Conveyor is that the parser implementation only depends on the syntax of the tool message generated by the model. Therefore, the number of parsers typically only depends on the number of models supported by the system, which is advantageous for tool developers.
@@ 4.1 Setup
We evaluate Conveyor on our testbed of servers with two Intel 10-core Xeon Gold 5215 CPUs (running at 2.5 GHz base frequency) and two NVIDIA GeForce RTX 3090 GPUs. We implement our system on top of PyTorch [21] with FlashInfer CUDA kernels [8]. Our Conveyor system is implemented in about 1.9K lines of Python. Our baseline is the same code but with tool partial execution disabled, where tool invocation always happens after decoding to the end-of-sequence (EOS) token.
@@
To the best of our knowledge, there are unfortunately no publicly available realistic datasets of tool-assisted LLM serving scenarios. Instead, we systematically investigate existing literature [14, 19, 12, 2, 27, 23, 15] and construct six scenarios for our evaluation.
@@
Due to the models’ capabilities limited by our testbed (the available GPU memory), we generate these synthetic workloads with our own prompts instead of using the original prompts of those papers. We use Mistral-7B-Instruct-v0.2 [20] for Python CodeGen and Planning workloads, and use Functionary-Small-v2.2 [10] for Search, Validation, Database, and Calculator workloads. We choose these two models because they can generate valid tool invocation instructions for us to run the evaluation except for the Validation workload. We pick Functionary for the Validation workload because we empirically found that it has a higher probability of generating format-correct requests. We set temperature to be 0, so every test for the same workload has the same LLM output. Performance variance across tests for the same workload is due to the performance variance in CPU/GPU processing and the Internet (for the Search and Planning workloads). Note that the effectiveness of Conveyor is not affected by the choice of LLM or prompts because the execution flow of LLM decoding instructions and invoking tools remains the same, which will be further elaborated in the following case study section.
@@ 4.3 Case Studies
For the CodeGen workload, we let Mistral-7B-Instruct-v0.2 [20] generates the Python code to plot a sine wave (Figure 3). Figure 4 shows the corresponding execution timelines with and without tool partial execution. Conveyor helps to reduce averagely 725 ms for the entire end-to-end serving latency, which is 3,918 ms on average without partial execution, leading to 26.3% improvement (see § 3.4) for this user request. It is worthwhile to note that a few import statements (e.g., line 2 and line 5) consume negligible time because Conveyor invokes the Python interpreter through fork, so the plotting script can reuse the modules already imported by Conveyor.
@@
For the Validation workload, we let Functionary-Small-v2.2 [10] generate a request for a local news service. The request is in JSON format, and the field “location” has to contain a valid city name and a state name. If the “location” only contains the city name, the request will be rejected by the service, and there is no point in sending such a request. Figure 7 shows the execution timeline. Without partial execution, such a check has to take place after the entire request is decoded. Partial execution allows the check to happen earlier and can abort immediately, saving the resources and the time for decoding subsequent tokens.